OpenStudy (anonymous):

how do u convert 27/40 into a percent

OpenStudy (jdoe0001):

\(\bf \cfrac{27}{40}\cdot 100\implies \square \%\)

OpenStudy (phi):

do you know how to divide 40 into 27 to get a decimal number ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

0.675

OpenStudy (phi):

to change a decimal to a percent, move the decimal point 2 places to the right (or the same thing, multiply the decimal by 100) what do you get ?

OpenStudy (anonymous):

67.5

OpenStudy (phi):

and add % to show it's a percent 67.5% yes, you got it.

OpenStudy (anonymous):

ok thank u

OpenStudy (phi):

some people learn a few percents by heart: 0.1 is 10% ½ is 0.5 is 50%

OpenStudy (mathstudent55):

A percent is simply a number divided by 100. What number, x, divided by 100 equals 27/40? \(\dfrac{27}{40} = \dfrac{x}{100}\) Cross multiply: 40x = 27 * 100 40x = 2700 Divide both sides by 40 x = 67.5 Answer is 67.5%

OpenStudy (anonymous):

27/40=.675 or 67.5%

OpenStudy (isaiah.feynman):